OpenStudy (anonymous):

If F(x)=x^2+4x, find the value(s) of x if the f(x) is 5.

OpenStudy (anonymous):

Put the x value inside the F(x) equation. F(5) = 5^2 + 4 * 5 F(5) = 25 + 20 F(5) = 45.

OpenStudy (anonymous):

ohh, the answer is 1,-5 though..

OpenStudy (anonymous):

how am I supposed to get it?

OpenStudy (anonymous):

but the answer is 1,-5 though..